If x is an acute angle and sin 2x = a find sin x + cos x
Answer by stanbon(75887)

If x is an acute angle and sin 2x = a find sin x + cos x
sin(2x)=2sinxcosx=a
sinxcosx=a/2
cosx=2/sinx
Therefore sinx + cosx = sinx+ 2/sinx
=(sinx)^2 + 2)/sinx
As with so much of trig your answer book answer may look
different than this.
Cheers,
Stan H.